Find the slope and y intercept of the line represented by each of the following equations. 2x-3y=6
Answer by checkley75(3666)

2X-3Y=6
-3Y=-2X+6
Y=-2X/-3+6/-3
Y=2X/3-2 THIS LINE HAS A SLOPE(m) OF 2/3 & A Y INTERCEPT(b) OF -2.